OpenStudy (jcmbourque19):

I have a few questions concerning logarithms... How can the properties of exponents help solve exponential equations? How do you solve 81^y=18? I know that its the same as log81 (18)=y but I don't know how to solve. Also I have logx-1 (8)=3 I really need explanations :/ I'm bad at math and this is so confusing.

OpenStudy (aravindg):

81^y=18 Taking log on both sides log 81^y= log 18 Now log a^n=n log a ylog 81=log 18 y=log 18/log 81 Do you understand?

OpenStudy (aravindg):

A similar method can be used to solve logx-1 (8)=3

OpenStudy (jcmbourque19):

I think so

OpenStudy (aravindg):

Good :)